OpenStudy (alison123):

solve: 504/224= 4x^2-10x-25/4x^2-12x+9 please help i started it but got so lost when solving it

OpenStudy (anonymous):

can you reduce the fraction on the left first?

OpenStudy (anonymous):

yeah it reduces a bunch start with \[\frac{9}{4}=\frac{4x^2-10x-25}{4x^2-12x+9}\]

OpenStudy (anonymous):

then cross multiply to get \[9(4x^2-12x+9)=4(4x^2-10x-25)\]

OpenStudy (anonymous):

hmm did i make a mistake? this has not solutions

OpenStudy (anonymous):

nope, i did not make a mistake, and it has no solutions

OpenStudy (alison123):

thanks anyways